OpenStudy (anonymous):

3x-y=0 2x-y=1 i am trying to solve these equations by graphing do you solve each equation first by putting it y=mx+b format

OpenStudy (anonymous):

suggestion, add y to both sides

OpenStudy (anonymous):

2x-y=1 2x-y+y=1+y 2x=1+y no we want to get rid of the 1 by subtracting 1 to both sides 2x-1=1-1+y 2x-1=y y=2x-1 that general process

OpenStudy (anonymous):

I don't understand how you would graph these equations

OpenStudy (anonymous):

the slope is m the y intercept is b y intercept means the point (0,b) slope is rise over run

OpenStudy (anonymous):

I am so confused on this graphing part how do you get the answer to know how to graph them

OpenStudy (anonymous):

the answer would be the point of intersection

OpenStudy (anonymous):

i understand that but i don't understand how get the answer to graph them

OpenStudy (anonymous):

did you graph them yet?

OpenStudy (anonymous):

no because I don't understand how to

OpenStudy (anonymous):

do you know what a coordinate point is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oy... thats a problem

OpenStudy (anonymous):

coordinate points come in the form of (x, y) 

OpenStudy (anonymous):

oh okay I know what you mean now

OpenStudy (anonymous):

teh easiest method to graph, cause i dont wnat to explain slope make a t chart  solve for y given a specific x, then plot the point then use a ruler to connect the point with a single line

OpenStudy (anonymous):

so the equation would 3x-y=0 3x-y+y=0 3x=0+y so my point would be 3,0 for the first equation

OpenStudy (anonymous):

no substitute for x, then solve for y if x =0, y=0 if x =1, y =3 if x=-1, y=-3

OpenStudy (anonymous):

in this case, b=0, b is the y intercept the y intercept comes in the form of (0,b) the y interecept is (0,0)

OpenStudy (anonymous):

slope means, from point to point, you go up 3 units and over 1